Calculate the power dissipated in the 2 ? resistor, if the voltage of the battarey is ?=24.1 V and the resistance of the resistor is R=7.2 Ohms? Express the answer to two decimal places.Calculate the power dissipated in the 2 ? r
Answer #1

1 .find the whole resistance of the circuit
[(5+1)x(2+1)]/[(5+1)+(2+1)]+7.2= 9.2 ohm

2. how much voltage going through the 2 ohm resister
(24.1 V/ 9.2 ohm)x2 ohm = 5.23V
[5.23 V/ (2+1)] x 2 ohm = 3.48 V

3. total power dissipate
P=(V`2)/R
P={(3.48)`2}/2
p= 6.0552 W